Quiz # 5 for Dr. Z.’s Number Theory

NAME: (print!) E-MAIL ADDRESS: (print!)

1. (5 points) Using divisiblity tests, determine whether 517286 is divisible by 11.

2. (5 points) What is the day of the week of Halloween 2100.